OpenStudy (anonymous):

The measure of a side of a square field is 2x^2+3What is its perimeter?

OpenStudy (anonymous):

2---------------sqaured (2x +3)

OpenStudy (anonymous):

(2x^2+3) = a^2 where a is the side of the square \[\sqrt{2x^2+3}=a\] perimeter = 4a perimeter = \[4\sqrt{2x^2+3}\]

OpenStudy (anonymous):

4 (2x^2 +3)

OpenStudy (anonymous):

oh sorry i thought 2x^2+3 is the area answer is \[4(2x^2+3)\]